Using the same information and data as was given in Problem 6, provide an 90% confidence interval for mean time advantage of right-hand over left-hand threads. Do you think that the time saved would be of practical importance if the task were performed many times - for example, by an assembly line worker? To help answer this question, find the mean time for right-hand threads as a percent of the mean time for left-hand threads. Be descriptive.

by the same information, the question is referring to the information/question below:

The design of controls and instruments has a large effect on how easily people can use them. A student project investigated this effect by asking 25 right-handed students to turn a knob (with their right hands) that moved an indicator by screw action. There were two identical instruments, one with a right-hand thread (the knob turns clockwise) and the other with a left-hand thread (the knob turns counter-clockwise). The table below gives the times required (in seconds) to move the indicator a fixed distance.

Subject ID

Right-hand Thread

Left-hand

Thread

1

113

137

2

105

105

3

130

133

4

101

108

5

138

115

6

118

170

7

87

103

8

116

145

9

75

78

10

96

107

11

122

84

12

103

148

13

116

147

14

107

87

15

118

166

16

103

146

17

111

123

18

104

135

19

111

112

20

89

93

21

78

76

22

100

116

23

89

78

24

85

101

25

88

123

  1. The project hoped to show that right-handed people find right-hand threads easier to use. State the appropriate Ho and Ha about the mean time required to complete the task.
  2. Carry out a test of your hypotheses. Give the P-value, report and explain your conclusion in a brief paragraph.
Answer #1

i.
TRADITIONAL METHOD
given that,
mean(x)=104.12
standard deviation , s.d1=15.796
number(n1)=25
y(mean)=117.44
standard deviation, s.d2 =27.262
number(n2)=25
I.
standard error = sqrt(s.d1^2/n1)+(s.d2^2/n2)
where,
sd1, sd2 = standard deviation of both
n1, n2 = sample size
standard error = sqrt((249.514/25)+(743.217/25))
= 6.302
II.
margin of error = t a/2 * (standard error)
where,
t a/2 = t -table value
level of significance, α = 0.1
from standard normal table,right tailed and
value of |t α| with min (n1-1, n2-1) i.e 24 d.f is 1.318
margin of error = 1.318 * 6.302
= 8.305
III.
CI = (x1-x2) ± margin of error
confidence interval = [ (104.12-117.44) ± 8.305 ]
= [-21.625 , -5.015]
-----------------------------------------------------------------------------------------------
DIRECT METHOD
given that,
mean(x)=104.12
standard deviation , s.d1=15.796
sample size, n1=25
y(mean)=117.44
standard deviation, s.d2 =27.262
sample size,n2 =25
CI = x1 - x2 ± t a/2 * Sqrt ( sd1 ^2 / n1 + sd2 ^2 /n2 )
where,
x1,x2 = mean of populations
sd1,sd2 = standard deviations
n1,n2 = size of both
a = 1 - (confidence Level/100)
ta/2 = t-table value
CI = confidence interval
CI = [( 104.12-117.44) ± t a/2 * sqrt((249.514/25)+(743.217/25)]
= [ (-13.32) ± t a/2 * 6.302]
= [-21.625 , -5.015]
-----------------------------------------------------------------------------------------------
interpretations:
1. we are 90% sure that the interval [-21.625 , -5.015] contains the true population proportion
2. If a large number of samples are collected, and a confidence interval is created
for each sample, 90% of these intervals will contains the true population proportion
90% confidence interval for mean time advantage of right-hand over left-hand threads
[-21.625 , -5.015]

ii.
Given that,
mean(x)=104.12
standard deviation , s.d1=15.796
number(n1)=25
y(mean)=117.44
standard deviation, s.d2 =27.262
number(n2)=25
null, Ho: u1 = u2
alternate, H1: u1 != u2
level of significance, α = 0.1
from standard normal table, two tailed t α/2 =1.711
since our test is two-tailed
reject Ho, if to < -1.711 OR if to > 1.711
we use test statistic (t) = (x-y)/sqrt(s.d1^2/n1)+(s.d2^2/n2)
to =104.12-117.44/sqrt((249.51362/25)+(743.21664/25))
to =-2.1138
| to | =2.1138
critical value
the value of |t α| with min (n1-1, n2-1) i.e 24 d.f is 1.711
we got |to| = 2.11377 & | t α | = 1.711
make decision
hence value of | to | > | t α| and here we reject Ho
p-value: two tailed ( double the one tail ) - Ha : ( p != -2.1138 ) = 0.045
hence value of p0.1 > 0.045,here we reject Ho
ANSWERS
---------------
a.
null, Ho: u1 = u2
alternate, H1: u1 != u2
test statistic: -2.1138
critical value: -1.711 , 1.711
decision: reject Ho
b.
p-value: 0.045
we have enough evidence to support the claim that There were two identical instruments, one with a right-hand thread (the knob turns clockwise) and the other with a left-hand thread (the knob turns counter-clockwise)